### v2

@@ Abstract
Alien species refer to non-native species introduced into an ecosystem, potentially causing harm to the environment, economy, or human health. Presence of confounding factors has so far prevented a comprehensive picture of relative importance of various drivers of such invasions. In this manuscript, we aim to develop and apply a general mixed additive Relational Event Model (REM) to describe the pattern of global invasions of alien species. An alien species invasion can be seen as a relational event, where the species - sender - reaches a region - receiver - at a specific time in history. We consider the First Record Database and focus on co-invasions by insects and plants between 1880 and 2005. REM aims to describe underlying hazard of each sender-receiver pair. Besides potentially time-varying, exogenous, and endogenous covariates, our mixed additive REM incorporates time-varying and random effects. Our efficient inference procedure relies on case-control sampling, yielding the same likelihood as that of a degenerate logistic regression. Resulting computational efficiency means that complex models for large dynamic networks can be estimated in seconds on a standard computer. We also present a framework for testing the goodness-of-fit of REMs via cumulative martingale-residuals. Implementation is performed through R package mgcv.
@@ 1 Introduction
Alien species refer to non-native species introduced into a new ecosystem, successfully overcoming challenges, such as geographical barriers and sustaining reproduction in the new location. Although not all alien species become invasive, this phenomenon is now widely recognised as a significant and widespread threat due to the resulting environmental damage and costs involved. The detrimental consequences of invasive species include reducing biodiversity, damaging ecosystems, and impacting human health (McNeely, 2001; Pyšek and Richardson, 2010).
Understanding the mechanisms driving the dispersal of alien species and evaluating the impact of various factors on their rate of spread are crucial for effectively addressing this issue. While some invasions occur through natural diffusion processes, many introductions are deliberately facilitated by humans for commercial or aesthetic purposes (Pyšek et al., 2020). Ecological, climatological, socioeconomic, historical, and geographical processes, along with species interactions, all contribute to this complex phenomenon.
Most of the aforementioned drivers are dynamic, changing over time and exhibiting significant fluctuations. For instance, international trade is widely acknowledged as one of the main factors driving biological invasions. Trade itself, and its influence on invasions, varies over time, as recent efforts such as international agreements and regular border surveillance aim to address the alien species invasions (Hulme, 2021).
@@
Several modelling approaches exist to describe these mechanisms. Species distribution models (SDMs), a broad family of models widely used in ecology, seek to explain presence and abundance of species as a response to environmental changes, identifying regions at risk of invasion (Bellard et al., 2016). These models have also been expanded to the community level by considering multiple species simultaneously. However, the adequacy of SDMs must be evaluated carefully, particularly regarding the spatial and temporal congruence between the variables (Araújo et al., 2019; A. Lee-Yaw et al., 2022). Given the complexity and global scale of the invasion process, realistic modelling requires a framework that can transparently incorporate multiple, alternative scientific hypotheses. Juozaitienė et al. (2023) developed the first formulations of a Relational Event Models (REMs) specific to the spread of alien species, considering both the time-varying nature of drivers, as well as of their effects. However, the computational cost of their statistical inference procedure necessitated simplifications in the model formulation. We propose a general mixed-effect additive relational event model that incorporates non-linear, time-varying, and random effects, combined with an efficient estimation approach. This implies that we can consider larger relational event networks consisting of multiple interacting taxonomic groups of species. Specifically, we are interested in studying the joint invasion patterns of vascular plants and insects.
In section Data Sources we describe the characteristics of the Alien Species First Record database and the additional sources of data that are employed in this study. Section Generative Species Invasion Model aims at defining the mixed additive relational event model involving time-varying covariates, time-varying effects and random effects. The following section proposes an efficient inference method for the model, as well as an associated goodness-of-fit technique. After showing the its accuracy in a target simulation study, we analyse the global joint invasion patterns of insects and vascular plants. We conclude with a brief evaluation of the method and insights gained on the alien species invasions, including some of its limits and potential future research paths.
@@ 2 Data Sources
In this section we describe the various data sources and records we consulted for our study. Our aim is to develop a parsimonious quasi-generative model for the spread of insects and vascular plants in recent history. First, we describe the information we have on the temporal establishment of certain species in specific regions. Then we focus on information relating to potential drivers of the invasion process. Additionally, a potential source of confounding is the existence of native ranges for the different species of plants and insects. The absence of invasion records would namely mistakenly suggest that the native region is unsuitable for that particular species, whereas the reverse is true. Therefore, we consider carefully what information is available on the native ranges of the various species of plants and insects.
@@ 2.2 Data on Potential Drivers of Alien Species Invasions
Natural invasion events over long geographical distances are relatively uncommon (Trakhtenbrot et al., 2005). In order to evaluate the precise role of distance, we consider the shortest distance among the regions in which the species is already present at that time. distance is computed referring to the closest borders, resulting in a value of zero for neighbouring regions. Source data for this driver consists of the R package geosphere (Hijmans et al., 2017). In order to avoid issues with outliers, distance is log-transformed.
In the existing literature, international trade has been acknowledged as a key factor for explaining the spread of alien species; furthermore, the value of import commodities is also referred to as a common proxy for the rate of alien species introductions (Seebens et al., 2018). Sometimes, the two terms are even used interchangeably (Hulme, 2021). Source information comes from Barbieri et al. (2009), and reports trade flows among countries. Trade values are not available for every pair of countries in each year considered. Juozaitienė et al. (2023) proposed an imputation method for the source data: when gaps emerge at the beginning of the observation period, they are replaced with zero; instead, gaps in intervening years of already recorded trade are imputed according to a log-linear model in case of growing trade. Our quantity of interest (trade) consists of the yearly commerce between already-invaded territories and the region of interest. Despite the use of the previously mentioned imputation method, this information may still not be computable from the data available. First, when pair of nations is never mentioned in the source data, the trade volume is presumed to be equal to zero, based on the assumption that if a value is never recorded, it is likely negligible. Second, when the couple has been mentioned but there is a non-imputed gap at the time of interest, the trade flow is evaluated at the latest year for which information is available. Finally, as trade varies by orders of magnitude, the values are log-transformed in the analysis.
Bellard et al. (2016) report the climate as a relevant driver for invasions. Our analysis relies on the inferred yearly values of near-surface air temperature from Watanabe et al. (2011). We introduce climatic dissimilarity for each country and species as the minimal temperature difference relative to all other countries where that species was already present before the year of interest.
@@
Colonial expansion, particularly European colonialism by the British, Spanish, Portuguese and Dutch, has been recognized as a factor in the dispersion of alien species (Dyer et al., 2017; Lenzner et al., 2022). Data on colonial ties are available from the COLDAT infrastructure (Becker, 2019), which reports the presence and the starting and ending date of colonial powers [^1] Belgium, Britain, France, Italy, Germany, Netherlands, Portugal, and Spain are the examined colonial powers. in the regions mentioned in Barbieri et al. (2009). In our analysis, each region either refers to the colonial power it belonged to or is classified as independent. For each species-region-year triplet, the covariate colonial ties is assigned a value of $1$ if the species is already present in the area of the colonial power that the invaded region belonged to, and 00 otherwise.
@@ 2.3 Native Range
A species’ native range (NR) is the collection of areas where it is indigenous. Slightly more liberally, we refer to NR as the set of sites where a species was already present before start of the analysis period, which in this context is 1880. This notion is relevant both in ecological terms and statistical terms. First, the ASFR database does not provide information on the origin of the species, but only the region it is invading. Knowledge of its NR allows us to identify which possible parts of the world this species hails from. Additionally, fascinating relationships between native and invaded environments may be examined (Hejda et al., 2015). Secondly, several of the factors mentioned in Section 2.2, in addition to being time-varying, are also endogenous — this means that these factors arise from the sequence of previously occurred invasions. For each species-region-year triplet, we need to know where the species was already present prior to that specific year. In this sense, the notion of NR is necessary statistically to compute the covariate values. Van Kleunen et al. (2019) and CABI Invasive Species Compendium (https://www.cabi.org, accessed 15.07.2016) are used as sources that describe the NRs of vascular plants and insects. Additionally, the ASFR invasions before 1880 were used to supplement the NRs.
@@ 3 Generative Species Invasion Model
An alien species invasion by species $s$ of a region $r$ in year $t$ can be considered a relational event. Relational event, involving a sender $s$ interacting with a receiver $r$ at time $t$, can be expressed as a triplet $e=(s,r,t)$. The FR invasion process (IP) can be modelled as a marked point process (MPP) $\{[t_{k},(s_{k},r_{k})];k\geq 1\}$, where species-region dyad can be considered a mark on event process. FR sequence can thus be expressed as,
@@
  • $\bm{b}$ includes random frailties, capturing additional heterogeneity.
@@ 3.1 Fixed, possibly Time-Varying, Effects
The covariate process $\bm{x}$ can consist of exogenous and endogenous variables. Exogenous covariates are external to the process, whereas endogenous covariates are functions of the IP itself. Due to the intricate nature of the IP process, all variables described in Table 1 are endogenous.
The impact of these time-varying drivers on the rate of occurrence may either be assumed to be fixed or allowed to vary over time. The nature of the effect for each variable is reported in Table 1. Time-varying effects for covariate $j$ are defined as $\bm{\beta}^{j}(t)=\bm{\beta}^{j^{\prime}}\bm{g}^{j}(t)$.
A possible option for $\bm{g}$ is a set of radial basis functions. The resulting thin plate splines are extremely flexible but involve a sizeable number of parameters. A thin plate regression spline is a low-rank approximation of thin plate splines that can be incorporated into a wide range of models (Wood, 2003). We consider the latter.
@@ 3.2 Random Effects
One of the goals of the study is to understand if heterogeneity of species invasiveness and region invasibility plays a significant role in the invasion process. Additionally, we want to explore the overall co-invasion patterns of species by examining if the presence of a species in a particular region affects the rate of invasion by other organisms in that region.
We thus include two kinds of random effects: monadic random intercepts capture heterogeneity of species and regions, whereas dyadic random intercepts model heterogeneity in the co-invasion of species pairs. It considers pairs consisting of the currently invading species and the most recent species to enter the area. Due to the difference in the order of magnitude of the number of insects and plants, we decided to model the species invasiveness random effect with between-strata heteroscedasticity.
@@ 4.1 Case-Control Partial Likelihood Inference via GAMs
Nested case-control (NCC) sampling (Borgan et al., 1995; Lerner and Lomi, 2020) considers for each event $(s,r)$ in $\mathcal{E}$ a reduced risk set composed of the event and $m-1$ non-events, sampled according to a given probability distribution $\pi_{t}(\cdot|sr)$. We consider the case in which $m=2$, i.e., at each time $t$ the sampled risk set $\bm{sr}$ consists of the event dyad $(s,r)$ and one non-event dyad $(s^{\ast},r^{\ast})$ randomly sampled from the complete risk set $\mathcal{R}(t)$, which consists of the pairs of species and regions that could be observed at time $t$.
@@ 4.2 Non-Parametric Estimation of the Baseline Hazard
The cumulative baseline function $\Lambda_{0}(t)=\int_{t_{0}}^{t}\lambda_{0}(u)\mathrm{d}u$ may be estimated non-parametrically. Borgan et al. (1995) and Borgan and Langholz (1997) propose an adaptation of the Breslow estimator for sampled cohort data. Consider the expected fitted linear predictors $\eta_{sr}(t)=\hat{\bm{\beta}}(t)^{\prime}\cdot\bm{x}_{sr}(t)+\mathbb{E}[\bm{b}\mid\mathcal{E}]^{\prime}\cdot\bm{z}_{sr}(t)=\hat{\bm{\gamma}}^{\prime}\bm{h}_{sr}(t)$, where $\hat{\bm{\gamma}}$ includes the GAMM fitted parameter and $\bm{h}_{sr}(t)$ the corresponding model matrix. Then, the Breslow estimator can be written as
$\widehat{\Lambda}_{0}(t|{\mathcal{H}}_{t^{-}})=\sum_{t_{k}\leq t}\dfrac{1}{\sum_{(s,r)\in SR_{t_{k}}}\exp{\left[\hat{\eta}{sr}(t_{k})\right]}\cdot\omega_{sr}(t_{k},SR_{t_{k}})}$ (8)
@@
Since we consider matched NCC with $m=2$ with a stratification $a$ of the species population into $a_{1}=$ plants and $a_{2}=$ insects, the weights in the baseline hazard estimate $\widehat{\Lambda}_{0a}$ simplify as $\omega_{sr}(t,\bm{sr})=|\mathcal{R}_{a}(t)|/2$.

### v3

@@ Abstract
Alien species refer to non-native species introduced by humans into an ecosystem, which can cause harm to the environment, economy, or human health. Although there is considerable literature on the subject, the presence of confounding factors has so far prevented a comprehensive picture of the relative importance of various drivers of such invasions. In this manuscript, we aim to develop and apply a general mixed additive relational event model to describe the pattern of global invasions of alien species. The diffusion of alien species can be regarded as a relational event, where the species – the sender – reaches a region – the receiver – at a specific time in history. We use the First Record Database, which contains all co-invasions by insects and plants between 1880 and 2005. A relational event model (REM) is employed to describe the underlying hazard of each species-region pair. Besides potentially time-varying, exogenous, and endogenous covariates, the mixed additive REM incorporates time-varying and random effects, allowing for taxa-specific baseline rates while accounting for the potential synergistic effect between plants and insects in the invasion process. Our efficient inference procedure relies on case-control sampling, yielding the same likelihood as that of a degenerate logistic regression. We propose fitting the mixed additive REM via a generalised additive model with random effects as 0-dimensional splines. The resulting computational efficiency means that complex models for large dynamic networks can be estimated in seconds on a standard computer. Furthermore, we present a framework for testing the goodness-of-fit of our mixed additive REM for the invasions by vascular plants and insects by means of cumulative martingale-residuals. Implementation is performed through the R package mgcv.
@@ 1 Introduction
Alien species refer to non-native species introduced by humans into a new ecosystem, successfully overcoming challenges, such as geographical barriers and sustaining reproduction in the new location. Although not all alien species become invasive, this phenomenon is now widely recognised as a significant and widespread threat due to the resulting environmental damage and costs involved. The detrimental consequences of invasive species include reducing biodiversity, damaging ecosystems, and impacting human health (McNeely, 2001; Pyšek and Richardson, 2010). Understanding the mechanisms driving the dispersal of alien species and evaluating the impact of various factors on their rate of spread are crucial for effectively addressing this issue. While some invasions occur through diffusion processes, many introductions are deliberately facilitated by humans for commercial or aesthetic purposes (Pyšek et al., 2020). Ecological, climatological, socioeconomic, historical, and geographical processes, along with species interactions, all contribute to this complex phenomenon. Most of the aforementioned drivers are dynamic, changing over time and exhibiting significant fluctuations. For instance, international trade is widely acknowledged as one of the main factors driving biological invasions. Trade itself, and its influence on invasions, varies over time, as recent efforts such as international agreements and regular border surveillance aim to address the alien species invasions (Hulme, 2021).
@@
Several modelling approaches exist to describe the species diffusion processes. Species distribution models (SDMs), a broad family of models widely used in ecology, seek to explain presence and abundance of species as a response to environmental changes, identifying regions at risk of invasion (Bellard et al., 2016). These models have also been expanded to the community level by considering multiple species simultaneously. However, the adequacy of SDMs must be evaluated carefully, particularly regarding the spatial and temporal congruence between the variables (Araújo et al., 2019; A. Lee-Yaw et al., 2022). Given the complexity and global scale of the invasion process, realistic modelling requires a framework that can transparently incorporate multiple, alternative scientific hypotheses. Juozaitienė et al. (2023) developed the first formulations of a relational event model (REM) specific to the spread of alien species, considering the first record as a relational event whereby a species is “connected” to the non-native region at the moment of invasion. This explanatory model describes the possibly time-varying effect of time-varying covariates on the occurrence of the invasion events. However, the computational cost of their statistical inference procedure necessitated simplifications in the model formulation. Instead, we propose a general mixed-effect additive relational event model that incorporates non-linear, time-varying, and random effects, combined with an efficient estimation approach. This implies that we can consider larger relational event networks consisting of multiple interacting taxonomic groups of species. Specifically, we are interested in studying the joint invasion patterns of vascular plants and insects. By modeling relational events involving species from different taxa jointly, we can estimate taxa-specific baseline rates while accounting for the influence of other taxa. Furthermore, this framework allows us to examine relationships among species across taxa, inferring how the species interactions influence the invasion process.
In section Data Sources we describe the characteristics of the Alien Species First Record database and the additional sources of data that are employed in this study. The section entitled Generative Species Invasion Model aims at defining the mixed additive relational event model involving time-varying covariates, time-varying effects and random effects. The following section proposes an efficient inference method for the model, as well as an associated goodness-of-fit technique. After showing its accuracy in a target simulation study, we analyse the global joint invasion patterns of insects and vascular plants. The code used in this study is publicly available on GitHub at https://github.com/martinaboschi/alienspecies.git. We conclude by discussing how our work relates to recent literature on invasion dynamics and species co-invasion structures, followed by a consideration of methodological limitations and potential future research paths.
@@ 2 Data Sources
In this section we describe the various data sources and records we consulted for our study. Our aim is to infer a parsimonious quasi-generative model for the spread of insects and vascular plants in recent history. Specifically, this model aims to describe the mechanisms of invasion using the available information, which is presented in this section. First, we describe the information we have on the temporal establishment of certain species in specific regions. Then we focus on information relating to potential drivers of the invasion process. Additionally, as a starting point of our analysis, we consider the native ranges at 1880. These native ranges are extremely important as all future invasions start from the current native ranges.
@@ 2.2 Data on Potential Drivers of Alien Species Invasions
[…]
trade $tr_{sr}(t)$ Logarithm of the sum of annual trade flows (in current United States dollars) between region $r$ and other countries that have been invaded by species $s$ before time $t$. This variable shows several missings: our imputation method is described in Section 2.2. Time-varying covariate with time-varying effect (Barbieri et al., 2009)
[…]
Invasion events over long geographical distances are relatively uncommon (Trakhtenbrot et al., 2005). In order to evaluate the precise role of distance, we consider the shortest distance among the regions in which the species is already present at that time. distance is computed referring to the closest borders, resulting in a value of zero for neighbouring regions. Source data for this driver consists of the R package geosphere (Hijmans et al., 2017). In order to avoid issues with outliers, distances are log-transformed.
In the existing literature, international trade has been acknowledged as a key factor for explaining the spread of alien species; the value of import commodities is a common proxy for the rate of alien species introductions (Seebens et al., 2018). Sometimes, the two terms are even used interchangeably (Hulme, 2021). Source information comes from Barbieri et al. (2009), and reports trade flows among countries. Trade values are not available for every pair of countries in each year considered. Juozaitienė et al. (2023) proposed an imputation method for the source data: when gaps emerge at the beginning of the observation period, they are replaced with zero; instead, gaps in intervening years of already recorded trade are imputed according to a log-linear model in case of growing trade. If a gap occurs at the end of the observation period, then the last available trade is carried forward. We define the quantity of interest, trade$(s,r)$, as the yearly commerce between already-invaded territories by $s$ and the region $r$ (recorded as bilateral trade flow in current United States dollars). As trade varies by orders of magnitude, the values are log-transformed in the analysis.
Bellard et al. (2016) report the climate as a relevant driver for invasions. Our analysis relies on the inferred yearly values of near-surface air temperature from Watanabe et al. (2011). Various studies have highlighted the significance of average annual temperature as a key component of climatic conditions (Seebens et al., 2018; Colling et al., 2025; Barni et al., 2012). While precipitation patterns exhibit greater variability than temperature under climate change (Finch et al., 2021), temperature remains strongly correlated with precipitation ($r=0.8$) (Colling et al., 2025). Furthermore, Barni et al. (2012) reported significant multicollinearity among climate variables and identified minimum temperature, mean temperature, and the heat sum for the spring season as the most important climatic factors. We introduce climatic dissimilarity for each country and species as the minimal temperature difference relative to all other countries where that species was already present before the year of interest.
@@
Colonial expansion, particularly European colonialism by the British, Spanish, Portuguese and Dutch, has been recognized as a factor in the dispersion of alien species (Dyer et al., 2017; Lenzner et al., 2022). Data on colonial ties are available from the COLDAT infrastructure (Becker, 2019), which reports the presence and the starting and ending date of colonial powers [^1] Belgium, Britain, France, Italy, Germany, Netherlands, Portugal, and Spain are the reported colonial powers. in the regions mentioned in Barbieri et al. (2009). In our analysis, each region either refers to the colonial power it belonged to or is classified as independent. For each species-region-year triplet, the covariate colonial ties is an indicator function of whether the species is already present in the area associated with the colonial power to which the region belongs.
@@ 2.3 Native Range
A species’ native range (NR) is the collection of areas where it is indigenous. Slightly more liberally, we refer to NR as the set of sites where a species was already present before start of the analysis period, which in this context is 1880. This notion is relevant both in ecological terms and statistical terms. First, the ASFR database does not provide information on the origin of the species, but only the region it is invading. Knowledge of its NR allows us to identify which possible parts of the world this species hails from. Additionally, fascinating relationships between native and invaded environments may be examined (Hejda et al., 2015). Secondly, several of the factors mentioned in the previous section are endogenous — this means that these factors arise from the sequence of previously occurred invasions. For each species-region-year triplet, we need to know where the species was already present prior to that specific year. In this sense, the notion of NR is necessary statistically to compute the covariate values. Van Kleunen et al. (2019) and CABI Invasive Species Compendium (https://www.cabi.org, accessed 15.07.2016) are used as sources that describe the NRs of vascular plants and insects. Additionally, the ASFR invasions before 1880 were used to supplement the NRs.
@@ 3 Generative Species Invasion Model
An alien species invasion by species $s$ of a region $r$ in year $t$ can be considered a relational event. A relational event, involving a sender $s$ interacting with a receiver $r$ at time $t$, can be expressed as a triplet $e=(s,r,t)$. The FR invasion process (IP) can be modelled as a marked point process (MPP) $\{[t_{k},(s_{k},r_{k})];k\geq 1\}$, where species-region dyad can be considered a mark on the event process. FR sequences can thus be expressed as,
@@
  • $\bm{b}$ are random frailties, capturing additional heterogeneity.
@@ 3.1 Fixed, possibly Time-Varying, Effects
The covariate process $\bm{x}$ can consist of exogenous and endogenous variables. Exogenous covariates are external to the process, whereas endogenous covariates are functions of the IP itself. Due to the intricate nature of the IP process, all variables described in Table 1 are endogenous. The impact of these time-varying drivers on the rate of occurrence may either be assumed to be fixed or allowed to vary over time. The nature of the effect for each variable is reported in Table 1. Time-varying effects for covariate $j$ are defined as $\bm{\beta}^{j}(t)=\sum_{i}\bm{\beta}^{j}_{i}\bm{g}^{j}_{i}(t)$. A possible option for $\bm{g}$ is a set of radial basis functions. The resulting thin plate splines are extremely flexible but involve a sizeable number of parameters. A thin plate regression spline is a low-rank approximation of thin plate splines that can be incorporated into a wide range of models (Wood, 2003). We consider the latter.
@@ 3.2 Random Effects
One of the goals of the study is to understand if heterogeneity of species invasiveness and region invasibility plays a significant role in the invasion process. Additionally, we want to explore the overall co-invasion patterns of species by examining if the presence of a species in a particular region affects the rate of invasion by other organisms in that region. We thus include two kinds of random effects: monadic random intercepts capture heterogeneity of species and regions, whereas dyadic random intercepts model heterogeneity in the co-invasion of species pairs. It considers pairs consisting of the currently invading species and the most recent species to enter the area. Due to the difference in the order of magnitude of the number of insects and plants, we decided to model the species invasiveness random effect with between-strata heteroscedasticity.
@@ 4.1 Case-Control Partial Likelihood Inference via GAMs
Nested case-control (NCC) sampling (Borgan et al., 1995; Lerner and Lomi, 2020) considers, for each event $(s,r)$ in $\mathcal{E}$, a reduced risk set composed of the event and $m-1$ non-events, sampled according to a given probability distribution $\pi_{t}(\cdot|sr)$. We consider the case in which $m=2$, i.e., at each time $t$ the sampled risk set $\bm{sr}$ consists of the event dyad $(s,r)$ and one non-event dyad $(s^{\ast},r^{\ast})$ randomly sampled from the complete risk set $\mathcal{R}(t)$, which consists of the pairs of species and regions that could be observed at time $t$.
@@ 4.2 Non-Parametric Estimation of the Baseline Hazard
The cumulative baseline function $\Lambda_{0}(t)=\int_{t_{0}}^{t}\lambda_{0}(u)\mathrm{d}u$ may be estimated non-parametrically. Borgan et al. (1995) and Borgan and Langholz (1997) propose an adaptation of the Breslow estimator for sampled cohort data. Consider the expected fitted linear predictors $\hat{\eta}_{sr}(t)=\hat{\bm{\beta}}(t)^{\prime}\cdot\bm{x}_{sr}(t)+\mathbb{E}[\bm{b}\mid\mathcal{E}]^{\prime}\cdot\bm{z}_{sr}(t)=\hat{\bm{\gamma}}^{\prime}\bm{h}_{sr}(t)$, where $\hat{\bm{\gamma}}$ includes the GAMM fitted parameter and $\bm{h}_{sr}(t)$ the corresponding model matrix. Then, the Breslow estimator can be written as
$\widehat{\Lambda}_{0}(t|{\mathcal{H}}_{t^{-}})=\sum_{t_{k}\leq t}\dfrac{1}{\sum_{(s,r)\in SR_{t_{k}}}\exp{\left[\hat{\eta}_{sr}(t_{k})\right]}\cdot\omega_{sr}(t_{k},SR_{t_{k}})}$ (8)
@@
Since we consider matched NCC with $m=2$ with a stratification $a$ of the species population into $a_{1}=$ plants and $a_{2}=$ insects, the weights in the baseline hazard estimate $\widehat{\Lambda}_{0a}$ simplify as $\omega_{sr}(t,\bm{sr})=|\mathcal{R}_{a}(t)|/2$. Alternatively, the baseline hazard in this setting may be estimated using shifted nested-case control sampling, as described in Lembo et al. (2024).